# Hidden yet quantifiable: A lower bound for confounding strength using randomized trials

### v2

@@ 1.1 Related work
Our approach is closely related to a line of work that proposes statistical tests for the presence of unobserved confounding. In particular, several works leverage randomized trials to detect unobserved confounding. These tests check for significant differences between average treatment effect estimates obtained from randomized and observational data [60, 64, 41]. More sophisticated approaches also test for differences in the group-level and conditional average treatment effect estimates [30, 29].
@@ 3.1 Statistical tests for $H_{0}(\Gamma)$
(i) If the estimators of the CATE sensitivity analysis bounds satisfy
$\|\mu_{\Gamma}^{\pm}-\hat{\mu}_{\Gamma}^{\pm}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=o_{\mathbb{P}}(n_{\operatorname{rct}}^{-1/2}),$
$\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)$ is a valid asymptotic test at level $\alpha$. (ii) If $\hat{\mu}_{\Gamma}^{+}$ and $\hat{\mu}_{\Gamma}^{-}$ are consistent estimators of the ATE sensitivity analysis bounds that satisfy
@@ 3.2 A lower bound on unobserved confounding strength
$\hat{\Gamma}_{\text{LB}}=\inf_{\Gamma}\{\Gamma:\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)=0\},$ (2)
that is, in words, the smallest $\Gamma$ such that the test accepts the null hypothesis. In practice, we compute $\hat{\Gamma}_{\text{LB}}$ with a grid search over values of $\Gamma$ starting from $1$ until the first test acceptance.
We show in the following proposition that $\hat{\Gamma}_{\text{LB}}$ is a valid lower bound for $\Gamma^{\star}$.
@@
[…]
@@
Note that by definition of $\hat{\Gamma}_{\text{LB}}$, we have that
$\mathbb{P}(\hat{\Gamma}_{\text{LB}}>\Gamma^{\star})$ $=\mathbb{P}(\cap_{\Gamma\leq\Gamma^{\star}}\{\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)=1\})$
$\leq\mathbb{P}(\hat{\phi}_{\mathrm{rct}}(\Gamma^{\star},\alpha)=1)\leq\alpha+o(1),$
@@ 4 Synthetic Experiments
First, we postulate that, for a fixed $\Gamma^{\star}$, the tightness of the lower bound $\hat{\Gamma}_{\text{LB}}$ improves when the confounder $U$ is more informative about the potential outcomes $(Y(1),Y(0))$. In our experiments, we choose the correlation between the unobserved confounder and one of the potential outcomes as a proxy measure of information,
@@
Intuitively, the sensitivity analysis bounds are tight for a specific $\Gamma^{\star}$ when $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}$ leads to a marginal distribution $\mathbb{P}^{\operatorname{os}}$ that maximally biases the estimable ATE. This situation occurs, for instance, when patients experiencing smaller outcomes are assigned to the control group while those with larger outcomes are in the treatment group. In this case, the sensitivity analysis bounds must be sufficiently large to include the true ATE and remain valid. Such a scenario is only possible if $U$ is very informative of $Y(1)$, captured by a high correlation coefficient. Conversely, when $Y(1)$ is independent of $U$, the true ATE is unaffected by the unobserved confounding, and the sensitivity bounds are unnecessarily conservative, leading to low power of the test and hence looser $\hat{\Gamma}_{\text{LB}}$. More formally, in Section A.3 we show that when the confounder $U$ is equal to $(Y(1),Y(0))$, the correlation coefficient $\rho_{u,y}=1$ and $\hat{\Gamma}_{\text{LB}}$ converges to $\Gamma^{\star}$ in the infinite sample limit. In contrast, when $U$ is independent of $(Y(1),Y(0))$, $\rho_{u,y}=0$ and $\hat{\Gamma}_{\text{LB}}=1$.
@@ 4.1 Datasets
Figure 3: For all the plots: the significance level is $\alpha=0.05$, $\phi^{\star}$ denotes the oracle test which rejects for $\Gamma<\Gamma^{\star}$, $\hat{\Gamma}_{\text{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\text{LB}}^{\widetilde{\mathrm{os}}}$ denote which test is used to compute $\hat{\Gamma}_{\text{LB}}$. First row with synthetic experiment choosing $\Gamma^{\star}=5$: Probability of rejection for different $\Gamma$ and average $\hat{\Gamma}_{\text{LB}}$ for the test for (a) small sample size: $n_{\operatorname{rct}}=2K,n_{\operatorname{os}}=2K$ and (b) large sample size: $n_{\operatorname{rct}}=20K,n_{\operatorname{os}}=20K$. $\hat{\Gamma}_{\text{LB}}$ for (c) increasing sample size of the observational study with $n_{\operatorname{rct}}=20K$ and (d) increasing correlation coefficient; $n_{\operatorname{rct}}=20K,n_{\operatorname{os}}=20K$. Second row with the semi-synthetic Hillstrom dataset choosing $\Gamma^{\star}=5$ and using “history” as unobserved confounder (except in (h)): Probability of rejection for different $\Gamma$ and average $\hat{\Gamma}_{\text{LB}}$ for (e) small sample size: $n_{\operatorname{rct}}=2300$, $n_{\operatorname{os}}=6150$ and (f) large sample size: $n_{\operatorname{rct}}=7680$, $n_{\operatorname{os}}=20500$. $\hat{\Gamma}_{\text{LB}}$ for (g) increasing $n_{\operatorname{os}}$ with $n_{\operatorname{rct}}=7680$ and (h) increasing correlation coefficient.
@@ 4.2 Experimental results
First, we observe in Figures 2(a)-2(b) and Figures 2(e)-2(f) that our tests are valid in all settings, i.e. they do not reject for strengths larger than $\Gamma^{\star}$. However, the statistical power substantially improves in the large sample size regime. In general, the performance of both tests aligns. In Figures 2(c) and 2(g), the lower bounds $\hat{\Gamma}_{\text{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\text{LB}}^{\widetilde{\mathrm{os}}}$ vary with the sample size of the observational study. We confirm that the $\hat{\phi}_{\widetilde{\mathrm{os}}}$ derives greater benefits from a larger observational study sample size than $\hat{\phi}_{\mathrm{rct}}$, as discussed in Section 3.1.
@@
Note that the tests in  Figure 2(a)-2(b) and  Figure 2(e)-2(f) are somewhat conservative: The probability of rejection for $\Gamma$ close to $\Gamma^{\star}$ is small, which leads to a rather loose lower bound estimate $\hat{\Gamma}_{\text{LB}}$. This is due to a fundamental limitation of the marginal sensitivity model that cannot be overcome without additional assumptions on how U affects Y, as discussed in Section A.3. We study here the effect of increasing the outcome-confounder correlation (Equation 3). Specifically, we generate observational datasets with a constant $\Gamma^{\star}$ but varying $\rho_{u,y}$, and report $\hat{\Gamma}_{\text{LB}}$ for both tests. For the synthetic experiments in Figure 2(d), we plot $\hat{\Gamma}_{\text{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\text{LB}}^{\widetilde{\mathrm{os}}}$ for $n=50$ distinct values of $\sigma^{2}_{Y(1)}\sim\operatorname{Unif}[0,1]$. For the semi-synthetic experiments in Figure 2(h), we depict $\hat{\Gamma}_{\text{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\text{LB}}^{\widetilde{\mathrm{os}}}$ for different hidden confounders $U$. Both plots confirm our hypothesis that higher $\rho_{u,y}$ correlates with a tighter lower bound $\hat{\Gamma}_{\text{LB}}$.
@@ 5 Real-world experiments
Linking back to the pipeline in Figure 1, we demonstrate how epidemiologists can use the lower bound $\hat{\Gamma}_{\text{LB}}$ to successfully differentiate between studies with significant confounding and those with negligible confounding. Specifically, we propose comparing $\hat{\Gamma}_{\text{LB}}$ with a critical value of $\Gamma$, estimated from the available observational data
$\hat{\Gamma}_{\operatorname{CT}}:=\inf\{\Gamma:0\in[\hat{\mu}^{-}_{\Gamma},\hat{\mu}^{+}_{\Gamma}]\}.$
In essence, $\hat{\Gamma}_{\operatorname{CT}}$ represents the minimum strength for which sensitivity analysis includes both positive and negative values of treatment effect, thereby invalidating the study results. Similar critical values have been proposed in the literature to assess the robustness of conclusions drawn from observational data, see e.g. VanderWeele and Ding [58], Jin et al. [32]. The most appropriate choice for the specific context should be determined by epidemiologists.
We flag an observational study as confounded if $\hat{\Gamma}_{\text{LB}}$ exceeds the critical value, i.e.
$\psi_{\mathrm{sens}}:=\mathbb{I}\{\hat{\Gamma}_{\text{LB}}>\hat{\Gamma}_{\operatorname{CT}}\}.$ (4)
@@
$\psi_{\mathrm{bin}}=\mathbb{I}\{\hat{\Gamma}_{\text{LB}}>1\}.$
@@
We consider two binary-valued outcomes: the presence of stroke and coronary heart disease within the follow-up period. We apply our procedure from Equation 4 to both the original dataset, which includes all patients (i.e. $t\leq 20$), and a subsampled dataset that only includes patients who were not previous users of HRT (i.e. $t=0$). Since the WHI study satisfies the criteria for a nested trial design, we calculate $\hat{\Gamma}_{\text{LB}}$ using our testing procedure $\hat{\phi}_{\widetilde{\mathrm{os}}}$. See Section B.3 for experimental details.
@@
$\hat{\Gamma}_{\mathrm{CT}}$ 1.017 1.172 1.017 1.164
$\hat{\Gamma}_{\mathrm{LB}}$ 1.052 1.207 1.009 1.224
$\psi_{\mathrm{bin}}$ 1 1 1 1
$\psi_{\mathrm{sens}}$ 1 1 0 1
@@ Results
For coronary heart disease, both algorithms flag the study as confounded when strong unobserved confounding is present ($t\leq 20$). However, when minimal unobserved confounding is present ($t=0$), our test does not flag the study, while $\psi_{\mathrm{bin}}$ does. This difference underscores our test’s capability to distinguish between small and large unobserved confounding, thereby addressing a limitation in the flagging procedures based on existing testing methods.
In the case of stroke, both $\psi_{\mathrm{sens}}$ and $\psi_{\mathrm{bin}}$ correctly flag the observational study, even when we adjust for the time since the start of treatment ($t=0$). This finding aligns with experts suggesting that additional unobserved confounding factors for stroke are still present after controlling for the time since the start of hormone replacement therapy [48].
@@ A.1 Remaining proofs
Hence, if $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{-}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o(1)$ and $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o(1)$, the theorem follows from the union bound. For brevity, we only prove $\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}$ as the proof for $\hat{T}_{\Gamma}^{-}$ is analogous.
@@
Further, let $n=n_{\operatorname{rct}}+n_{\operatorname{os}}$, where $n_{\operatorname{rct}}\ll n_{\operatorname{os}}$, and define
@@
$\sqrt{n_{\operatorname{rct}}}\left(\bar{Z}_{n}-\mu\right)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma),$
and it follows from the Cramér-Wold theorem that
$\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right).$
We now add and subtract the estimator of the CATE bounds
$\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)$ $=\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\hat{\mu}_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)+\sqrt{n_{\operatorname{rct}}}\left(\underbrace{\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})}_{O_{P}(\frac{1}{\sqrt{n_{\operatorname{os}}}})}\right)$
$=\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\hat{\mu}_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)+o_{P}(1),$
where in the last equality we have used $\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=o_{P}(n_{\operatorname{rct}}^{-1/2})$ and $n_{\operatorname{rct}}\ll n_{\operatorname{os}}$. Hence, we also have
$\sqrt{n_{\operatorname{rct}}}\left(\hat{\mu}_{\Gamma}^{+}-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right).$ (6)
@@
$\lim_{n_{\operatorname{rct}}\to\infty}P_{H_{0}}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $=\lim_{n_{\operatorname{rct}}\to\infty}P_{H_{0}}\left(\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}+2\hat{\sigma}^{+}_{\Gamma}\hat{\sigma}}}\leq z_{\alpha/2}\right)$
$=\lim_{n_{\operatorname{rct}}\to\infty}P_{H_{0}}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})}{\sqrt{(\sigma^{+}_{\Gamma})^{2}+\sigma^{2}+2\sigma^{+}_{\Gamma}\sigma}}\leq z_{\alpha/2}\right)$
$\leq\lim_{n_{\operatorname{rct}}\to\infty}P_{H_{0}}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})}{\sqrt{(\sigma^{+}_{\Gamma})^{2}+\sigma^{2}-2\xi_{\Gamma}}}\leq z_{\alpha/2}\right)$

### v3

@@ 1.1 Related work
Our approach is closely related to a line of work that proposes statistical tests for the presence of unobserved confounding. In particular, several works leverage randomized trials to detect unobserved confounding. These tests check for significant differences between average treatment effect estimates obtained from randomized and observational data [61, 65, 42, 30]. More sophisticated approaches also test for differences in conditional average treatment effect estimates [31] and account for right-censored outcomes [15].
@@ 3.1 Statistical tests for $H_{0}(\Gamma)$
(i) If it holds that $\underset{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}{\lim}n_{\operatorname{rct}}/n_{\operatorname{os}}=0$ and the estimators of the CATE sensitivity analysis bounds satisfy
$\|\mu_{\Gamma}^{\pm}-\hat{\mu}_{\Gamma}^{\pm}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=O_{\mathbb{P}^{\operatorname{os}}}(n_{\operatorname{os}}^{-1/2}),$
$\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)$ is a valid asymptotic test at level $\alpha$.
(ii) If $\hat{\mu}_{\Gamma}^{+}$ and $\hat{\mu}_{\Gamma}^{-}$ are consistent estimators of the ATE sensitivity analysis bounds that satisfy
@@ 3.2 A lower bound on unobserved confounding strength
$\hat{\Gamma}_{\texttt{LB}}=\inf_{\Gamma}\{\Gamma:\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)=0\},$ (2)
that is, in words, the smallest $\Gamma$ such that the test accepts the null hypothesis. In practice, we compute $\hat{\Gamma}_{\texttt{LB}}$ with a grid search over values of $\Gamma$ starting from $1$ until the first test acceptance.
We show in the following proposition that $\hat{\Gamma}_{\texttt{LB}}$ is a valid lower bound for $\Gamma^{\star}$.
@@
Let $\hat{\Gamma}_{\texttt{LB}}$ be as in Equation (2) for a fixed significance level $\alpha$. Then, under Assumptions 2.12.3 and the setting described in Section 2, $\hat{\Gamma}_{\texttt{LB}}$ is an asymptotically valid lower bound, i.e.
$\mathbb{P}(\hat{\Gamma}_{\texttt{LB}}\leq\Gamma^{\star})\geq 1-\alpha-o_{\mathbb{P}}(1).$
@@
Note that by definition of $\hat{\Gamma}_{\texttt{LB}}$, we have that
$\mathbb{P}(\hat{\Gamma}_{\texttt{LB}}>\Gamma^{\star})$ $=\mathbb{P}(\cap_{\Gamma\leq\Gamma^{\star}}\{\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)=1\})$
$\leq\mathbb{P}(\hat{\phi}_{\mathrm{rct}}(\Gamma^{\star},\alpha)=1)\leq\alpha+o_{\mathbb{P}}(1),$
@@ 4 Synthetic Experiments
First, we postulate that, for a fixed $\Gamma^{\star}$, the tightness of the lower bound $\hat{\Gamma}_{\texttt{LB}}$ improves when the confounder $U$ is more informative about the potential outcomes $(Y(1),Y(0))$. In our experiments, we choose the correlation between the unobserved confounder and one of the potential outcomes as a proxy measure of information,
@@
Intuitively, the sensitivity analysis bounds are tight for a specific $\Gamma^{\star}$ when $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}$ leads to a marginal distribution $\mathbb{P}^{\operatorname{os}}$ that maximally biases the estimable ATE. This situation occurs, for instance, when patients experiencing smaller outcomes are assigned to the control group while those with larger outcomes are in the treatment group. In this case, the sensitivity analysis bounds must be sufficiently large to include the true ATE and remain valid. Such a scenario is only possible if $U$ is very informative of $Y(1)$, captured by a high correlation coefficient. Conversely, when $Y(1)$ is independent of $U$, the true ATE is unaffected by the unobserved confounding, and the sensitivity bounds are unnecessarily conservative, leading to low power of the test and hence looser $\hat{\Gamma}_{\texttt{LB}}$. More formally, in Section A.3 we show that when the confounder $U$ is equal to $(Y(1),Y(0))$, the correlation coefficient $\rho_{u,y}=1$ and $\hat{\Gamma}_{\texttt{LB}}$ converges to $\Gamma^{\star}$ in the infinite sample limit. In contrast, when $U$ is independent of $(Y(1),Y(0))$, $\rho_{u,y}=0$ and $\hat{\Gamma}_{\texttt{LB}}=1$.
@@ 4.1 Datasets
Figure 3: For all the plots: the significance level is $\alpha=0.05$, $\phi^{\star}$ denotes the oracle test which rejects for $\Gamma<\Gamma^{\star}$, $\hat{\Gamma}_{\texttt{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\texttt{LB}}^{\widetilde{\mathrm{os}}}$ denote which test is used to compute $\hat{\Gamma}_{\texttt{LB}}$. First row with synthetic experiment choosing $\Gamma^{\star}=5$: Probability of rejection for different $\Gamma$ and average $\hat{\Gamma}_{\texttt{LB}}$ for the test for (a) small sample size: $n_{\operatorname{rct}}=2K,n_{\operatorname{os}}=2K$ and (b) large sample size: $n_{\operatorname{rct}}=20K,n_{\operatorname{os}}=20K$. $\hat{\Gamma}_{\texttt{LB}}$ for (c) increasing sample size of the observational study with $n_{\operatorname{rct}}=20K$ and (d) increasing correlation coefficient; $n_{\operatorname{rct}}=20K,n_{\operatorname{os}}=20K$. Second row with the semi-synthetic Hillstrom dataset choosing $\Gamma^{\star}=5$ and using “history” as unobserved confounder (except in (h)): Probability of rejection for different $\Gamma$ and average $\hat{\Gamma}_{\texttt{LB}}$ for (e) small sample size: $n_{\operatorname{rct}}=2300$, $n_{\operatorname{os}}=6150$ and (f) large sample size: $n_{\operatorname{rct}}=7680$, $n_{\operatorname{os}}=20500$. $\hat{\Gamma}_{\texttt{LB}}$ for (g) increasing $n_{\operatorname{os}}$ with $n_{\operatorname{rct}}=7680$ and (h) increasing correlation coefficient.
@@ 4.2 Experimental results
First, we observe in Figures 3(a)-3(b) and Figures 3(e)-3(f) that our tests are valid in all settings, i.e. they do not reject for strengths larger than $\Gamma^{\star}$. However, the statistical power substantially improves in the large sample size regime. In general, the performance of both tests aligns. In Figures 3(c) and 3(g), the lower bounds $\hat{\Gamma}_{\texttt{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\texttt{LB}}^{\widetilde{\mathrm{os}}}$ vary with the sample size of the observational study. We confirm that the $\hat{\phi}_{\widetilde{\mathrm{os}}}$ derives greater benefits from a larger observational study sample size than $\hat{\phi}_{\mathrm{rct}}$, as discussed in Section 3.1.
@@
Note that the tests in  Figure 3(a)-3(b) and  Figure 3(e)-3(f) are somewhat conservative: The probability of rejection for $\Gamma$ close to $\Gamma^{\star}$ is small, which leads to a rather loose lower bound estimate $\hat{\Gamma}_{\texttt{LB}}$. This is due to a fundamental limitation of the marginal sensitivity model that cannot be overcome without additional assumptions on how U affects Y, as discussed in Section A.3. We study here the effect of increasing the outcome-confounder correlation (Equation 3). Specifically, we generate observational datasets with a constant $\Gamma^{\star}$ but varying $\rho_{u,y}$, and report $\hat{\Gamma}_{\texttt{LB}}$ for both tests. For the synthetic experiments in Figure 3(d), we plot $\hat{\Gamma}_{\texttt{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\texttt{LB}}^{\widetilde{\mathrm{os}}}$ for $n=50$ distinct values of $\sigma^{2}_{Y(1)}\sim\operatorname{Unif}[0,1]$. For the semi-synthetic experiments in Figure 3(h), we depict $\hat{\Gamma}_{\texttt{LB}}^{\mathrm{rct}}$ and $\hat{\Gamma}_{\texttt{LB}}^{\widetilde{\mathrm{os}}}$ for different hidden confounders $U$. Both plots confirm our hypothesis that higher $\rho_{u,y}$ correlates with a tighter lower bound $\hat{\Gamma}_{\texttt{LB}}$.
@@ 5 Real-world experiments
Linking back to the pipeline in Figure 1, we demonstrate how epidemiologists can use the lower bound $\hat{\Gamma}_{\texttt{LB}}$ to successfully differentiate between studies with significant confounding and those with negligible confounding. Specifically, we propose comparing $\hat{\Gamma}_{\texttt{LB}}$ with a critical value of $\Gamma$, estimated from the available observational data
$\hat{\Gamma}_{\texttt{CT}}:=\inf\{\Gamma:0\in[\hat{\mu}^{-}_{\Gamma},\hat{\mu}^{+}_{\Gamma}]\}.$
In essence, $\hat{\Gamma}_{\texttt{CT}}$ represents the minimum strength for which sensitivity analysis includes both positive and negative values of treatment effect, thereby invalidating the study results. Similar critical values have been proposed in the literature to assess the robustness of conclusions drawn from observational data, see e.g. VanderWeele and Ding [59], Jin et al. [33]. The most appropriate choice for the specific context should be determined by epidemiologists.
We flag an observational study as confounded if $\hat{\Gamma}_{\texttt{LB}}$ exceeds the critical value, i.e.
$\psi_{\texttt{sensi}}:=\mathbb{I}\{\hat{\Gamma}_{\texttt{LB}}>\hat{\Gamma}_{\texttt{CT}}\}.$ (4)
@@
$\psi_{\texttt{bin}}=\mathbb{I}\{\hat{\Gamma}_{\texttt{LB}}>1\}.$
@@
We consider two binary-valued outcomes: the presence of stroke and coronary heart disease within the follow-up period. We apply our procedure from Equation 4 to both the original dataset, which includes all patients (i.e. $t\leq 20$), and a subsampled dataset that only includes patients who were not previous users of HRT (i.e. $t=0$). Since the WHI study satisfies the criteria for a nested trial design, we calculate $\hat{\Gamma}_{\texttt{LB}}$ using our testing procedure $\hat{\phi}_{\widetilde{\mathrm{os}}}$. See Section B.3 for experimental details.
@@
$\hat{\Gamma}_{\texttt{CT}}$ 1.017 1.172 1.017 1.164
$\hat{\Gamma}_{\texttt{LB}}$ 1.052 1.207 1.009 1.224
$\psi_{\texttt{bin}}$ 1 1 1 1
$\psi_{\texttt{sensi}}$ 1 1 0 1
@@ Results
For coronary heart disease, both algorithms flag the study as confounded when strong unobserved confounding is present ($t\leq 20$). However, when minimal unobserved confounding is present ($t=0$), our test does not flag the study, while $\psi_{\texttt{bin}}$ does. This difference underscores our test’s capability to distinguish between small and large unobserved confounding, thereby addressing a limitation in the flagging procedures based on existing testing methods.
In the case of stroke, both $\psi_{\texttt{sensi}}$ and $\psi_{\texttt{bin}}$ correctly flag the observational study, even when we adjust for the time since the start of treatment ($t=0$). This finding aligns with experts suggesting that additional unobserved confounding factors for stroke are still present after controlling for the time since the start of hormone replacement therapy [49].
@@ A.1 Remaining proofs
Hence, if $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{-}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o_{\mathbb{P}}(1)$ and $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o_{\mathbb{P}}(1)$, the theorem follows from the union bound. For brevity, we only prove $\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}$ as the proof for $\hat{T}_{\Gamma}^{-}$ is analogous.
@@
Further, we define
@@
$\sqrt{n_{\operatorname{rct}}}\left(\bar{Z}_{n}-\mu\right)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty.$
Thus, it follows from the Cramér-Wold theorem that
$\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty.$
It remains to show that asymptotic normality also holds when we use the empirical estimate $\hat{\mu}_{\Gamma}^{+}$. To do so, we prove the following convergence in probability
$\left|\frac{1}{\sqrt{n_{\operatorname{rct}}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right|=o_{\mathbb{P}^{\operatorname{os}}}(1),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty.$ (6)
Then, by Slutsky’s theorem, we have
$\sqrt{n_{\operatorname{rct}}}\left(\hat{\mu}_{\Gamma}^{+}-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty\;\;\text{and}\;\;n_{\operatorname{os}}\to\infty.$ (7)
First, we observe that the mean of the LHS in Equation 6 converges to zero in probability:
$\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[\frac{1}{\sqrt{n_{\operatorname{rct}}}}\left|\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right|\right]\leq\sqrt{n_{\operatorname{rct}}}~{}\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[\left|\mu_{\Gamma}^{+}(X)-\hat{\mu}_{\Gamma}^{+}(X)\right|\right]\leq\sqrt{n_{\operatorname{rct}}}~{}\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=o_{\mathbb{P}^{\operatorname{os}}}(1),$
where in the last equality we have used $\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=O_{\mathbb{P}^{\operatorname{os}}}(n_{\operatorname{os}}^{-1/2})$ and $n_{\operatorname{rct}}\ll n_{\operatorname{os}}$. Next, we show that the variance term also converges to zero in probability, i.e.
$\operatorname{Var}_{\mathbb{P}^{\operatorname{rct}}}\left[\frac{1}{\sqrt{n_{\operatorname{rct}}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right]\leq\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[\left(\mu_{\Gamma}^{+}(X)-\hat{\mu}_{\Gamma}^{+}(X)\right)^{2}\right]=\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}^{2}=o_{\mathbb{P}^{\operatorname{os}}}(1),$
and the statement in Equation 6 follows.
@@
$\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $=\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}+2\hat{\sigma}^{+}_{\Gamma}\hat{\sigma}}}\leq z_{\alpha/2}\right)$
$=\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})}{\sqrt{(\sigma^{+}_{\Gamma})^{2}+\sigma^{2}+2\sigma^{+}_{\Gamma}\sigma}}\leq z_{\alpha/2}\right)$
$\leq\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})}{\sqrt{(\sigma^{+}_{\Gamma})^{2}+\sigma^{2}-2\xi_{\Gamma}}}\leq z_{\alpha/2}\right),$